OpenStudy (anonymous):

what is the short way to solve (2+h)^3

OpenStudy (dumbcow):

https://en.wikipedia.org/wiki/Binomial_theorem

OpenStudy (anonymous):

Binomial theorem is awesome :)

OpenStudy (anonymous):

\[(a+b)^3 = (a+b)(a^2 + 2ab + b^2) = a^3 + 3a^2b + 3ab^2 + b^3\]

OpenStudy (anonymous):

the exponents of a start at 3 and go down: 3, 2, 1, 0 and the exponents of b go upwards: 0, 1, 2, 3